# RL Zero: Zero-Shot Language to Behaviors without any supervision

### v1

@@ 3 Preliminaries
Step 2 – Zero-Shot Inference: Once the successor features and basic features are obtained, the zero-shot policy can be obtained given any task by inferring the optimal $z$ that parametrizes the policy. If the task is specified using reward function $r$, the optimal $z$ parameterizing the policy can be obtained by:
$z^{*}=\operatorname*{arg\,max}_{z}\mathbb{E}_{\rho}[M^{\pi_{z}}r].$ (5)
For $M$ parameterized as Equation 3, and $\pi_{z}$ defined as Equation 4, $z^{*}$ can be found in closed form as $z^{*}=\mathbb{E}_{\rho}[\varphi(s)r(s)]$ (Touati & Ollivier, 2021). For other representations of successor measures, such as PSM (Agarwal et al., 2024), this $z^{*}$ can be found by solving a simple constraint optimization problem. Following prior work (Pirotta et al., 2023), we shall refer to this framework of pretraining successor features that can be used for zero-shot inference as Behavior Foundation Models (BFM).
Setting: The agent has access to an offline interaction dataset $d^{O}$ with density $\rho$ consisting of $\{s,o,a,s^{\prime}\}$ reward-free tuples where $o$ denotes the observation space (images for the purpose of this paper).
4 RLZero: Zero-shot prompt to policy
RLZero uses components trained with unsupervised learning to map language to behaviors. In the following sections, we describe the steps involved in detail: First, we present how an imagined trajectory is generated from a prompt. Then, we discuss how the imagined trajectory for a given prompt is projected to real observations of an agent. Finally, we describe the zero-shot procedure for inferring a policy that matches the behavior in the imagined trajectory.
4.1 Imagine: Generative Video Modeling
Grounding language to tasks in robotics has historically (Goyal et al., 2021a; Jang et al., 2022; O’Neill et al., 2023) required costly annotation labels that map language to task examples specified through image or state trajectories. Large video-language foundation models (ViFMs) help lift that requirement by training on vast amounts of internet videos, thus giving us a rich prior of grounding language commands to videos. A common issue in robotics, both simulated and real world, is that the domains often differ from real-world videos. We rely on a generative video modeling approach, GenRL (Mazzaglia et al., 2024), that uses a task encoder provided by an off-the-shelf ViFM (InternVideo (Wang et al., 2022)) along with a GRU architecture to imagine the video in latent space trained to be reconstructed to the environment domain. Training the video generation model does not require labels mapping language to tasks and is fully unsupervised. With an increase in ViFM scaling and developments in controllable video generation (Bruce et al., 2024; Hu et al., 2022; Ni et al., 2023; Chen et al., 2025), a few examples of environment domain may be sufficient to generate high quality in-domain imagination. Thus, given a language instruction $e^{l}$, we obtain a sequence of frames $(i_{1},i_{2}...i_{T})=V\!M(e^{l})$ that represents an imagination of what the task looks like in the environment domain. These imaginations are not expected to adhere to the agent’s environment dynamics. Figure 2 shows an example of what these imaginations look like using an off-the-shelf video generation model (Mazzaglia et al., 2024).
4.1.1 Grounding imagination in real-states
Algorithm 1 RLZero
2:  Given: text prompt $t$
The imaginings produced by VLMs can be noisy, unrealizable, and not exactly representative of the domain. We propose to use a similarity-based retrieval for the nearest frames in the dataset of the agent’s prior environmental interactions $d^{O}$ to project the imagined trajectories in real observations. This step allows us to match imagination to real observations in the semantic space, giving us the flexibility to use imaginations in differing domains than the agent (e.g. cross-embodiment discussed in Section 5). In this work, we use a performant image embedding approach for retrieval, SigLIP (Zhai et al., 2023), to map both the imagined frame and agent observation to a latent embedding space, which is trained for similarity matching with a contrastive objective. We use an encoding function $\mathcal{E}:\mathcal{I}\rightarrow\mathcal{Z}$ to individually map a sequence of images to shared text-image embedding space. For each consecutive $k$ length sequence of frames in the imagined trajectory, we output the following agent observations:
$o_{t-k:t}=\operatorname*{arg\,max}_{o_{t-k:t}}\frac{\mathcal{E}(o_{t-k:t})\cdot\mathcal{E}(i_{t-k:t})}{\|\mathcal{E}(o_{t-k:t})\|\|\mathcal{E}(i_{t-k:t})\|}~{}~{}~{}\forall t\in[T].$ (6)
Using $k$ previous frames allows us to identify state variables that correspond to quantities such as velocity, acceleration, etc, that are not identifiable from a single frame. This technique of ‘frame-stacking’ is commonly used in visual RL (Laskin et al., 2020) and has the interpretation of making the observation inputs Markov. Using the offline interaction dataset, we find corresponding proprioceptive states in addition to the real observation that we will subsequently use for distribution matching. While we rely on proprioceptive states in this work to solve distribution matching, in general our approach is not limited as BFMs may be trained with image observations if the state information is unavailable.
Figure 3: Grounding Imagination in Real Observations: We use nearest image retrieval defined by cosine similarity in the embedding space to output a real observation from the dataset that matches the imagined observation.
4.2 Imitate: Distribution Matching with Zero-Shot RL
The successor measure-based family of BFMs (Touati & Ollivier, 2021; Touati et al., 2023; Agarwal et al., 2024) captures the state visitation distribution of any policy in their learned bag of skills parameterized by $z$ and given by $\rho^{\pi_{z}}(s)=\mathbb{E}_{s_{0},a_{0}\sim\rho}[M^{\pi_{z}}(s_{0},a_{0},s)]$, where $M$ denotes the successor measure. We take the distribution matching perspective of imitation learning (Ghasemipour et al., 2020; Ni et al., 2021; Sikchi et al., 2024) and minimize the distance between the state visitation distributions of the grounded imagined trajectories (expert) and the policy denoted by $\rho^{E}$ and $\rho^{\pi_{z}}$ respectively:
$z_{imit}=\operatorname*{arg\,min}_{z}\text{distance}(\rho^{\pi_{z}}(s),\rho^{E}(s)),$ (7)
where the distance can be chosen to be mean-squared error, $f$-divergence, Integral Probability Metrics (IPM), etc. In general, minimizing the distance via gradient descent can provide a solution $z_{imit}$ to distribution matching. For the special case of KL divergence, Theorem 1 shows that $z_{imit}$ can be obtained in closed form using a learned distribution ratio between expert and offline interaction dataset $\rho^{E}/\rho$.
Theorem 1.
Define $J(\pi,r)$ to be the expected return of a policy $\pi$ under reward $r$. For an offline dataset $d^{O}$ with density $\rho$, a learned log distribution ratio: $\nu(s)=\log(\frac{\rho^{E}(s)}{\rho(s)})$, $D_{KL}(\rho^{\pi},\rho^{E})\leq-J(\pi,r^{imit})+D_{KL}(\rho^{\pi}(s,a),\rho(s,a))$ where $r^{imit}(s)=\nu(s)~{}\forall s$. The corresponding $z_{imit}$ minimizing the upper bound is given by $z_{imit}=\mathbb{E}_{\rho}[{r^{imit}(s)}\varphi(s)]=\mathbb{E}_{\rho^{E}}[\frac{\nu(s)}{e^{\nu(s)}}\varphi(s)]$ where $\varphi$ denoted state features learned by the BFM.
Thus, with the reward functions specified by $r^{imit}$, we can use the closed form solution of $z_{imit}=\mathbb{E}_{\rho}{[\varphi(s)r^{imit}(s)]}$ to retrieve the policy that mimics the grounded imagined behavior. This reward function requires learning a discriminator to obtain the distribution ratio, which can lead to instabilities, but a heuristic yet performant alternative is to use a shaped reward function $r(s)=e^{\nu(s)}$, similar to Pirotta et al. (2023), which allows zero-shot inference ($z_{imit}=\mathbb{E}_{\rho^{E}}{[\varphi(s)]}$) without learning a discriminator. We compare both approaches in Appendix C.1. The performance for both these methods are almost identical and we defer to the latter one in all our experiments. Using a state-only visitation matching objective can be limiting in the case where environmental dynamics permit the permutation of observation sequences that result in the same visitation distribution. This limitation can be relaxed by instead matching visitation on $\{s,s^{\prime}\}$. This requires minimal changes to training the BFM, but we found this to not be a limitation with the environments we consider.
5 Experiments
Our experiments seek to understand the quality of behaviors that the RLZero approach is able to produce given language prompts. The evaluation of these behaviors can be challenging as, unlike the traditional RL setting, we do not have access to a ground truth reward function. Instead, we have prompts that can be inherently ambiguous but reflect the reality of human-robot interaction. An obvious evaluation metric is to ask humans how much the generated behavior resembles their expectation of the behavior given the prompt. We use multimodal LLMs to evaluate such preferences as a proxy to human preferences, as recent studies (Chen et al., 2024) have shown them to be correlated (up to 79.3%).
Setup: We consider four DMC control tasks (Cheetah, Walker, Quadruped, and Stickman (Mazzaglia et al., 2024)). The Stickman environment reflects a human morphology with challenging control due to a large observation and action space. For task-conditioned video-generation we use off-the-shelf models from Mazzaglia et al. (2024). To obtain the nearest observation corresponding to the imagined image, we use SigLIP (Zhai et al., 2023), a state-of-the-art image-text embedding model. In all environments except Stickman, we collect data using RND (Burda et al., 2018) using the protocol specified in ExoRL (Yarats et al., 2022). For Stickman, we augment our dataset with replay buffers of the agent trained for run and walk behaviors, as obtaining meaningful tasks with pure random exploration is difficult with the large action-space of Stickman. The detailed composition of the datasets can be found in Appendix B.3. The behavior foundation model can be trained using any zero-shot RL method using successor features (Park et al., 2024a; Touati et al., 2023; Agarwal et al., 2024). In our experiments, we use the Forward-Backward zero-shot RL algorithm (Touati et al., 2023) using offline datasets.
Setting and Baselines: For our evaluations, we consider the setting where the agent has no access to the simulator during test time. This setting truly reflects the ability of the agents to use prior exploratory data to learn meaningful behaviors. We compare state-of-the-art model-free offline RL algorithms that are capable of learning from purely offline data. For RL algorithms, the reward is obtained as suggested by recent approaches that show the utility of VLM’s as reward functions (Baumli et al., 2023; Rocamonde et al., 2023). We consider two sources of reward: Image-language cosine similarity using SigLIP embedding, and Video-language cosine similarity using InternVideo2 embeddings. Video language embeddings take into account context and can potentially lead to more accurate reward estimation. Once the rewards are available, we use TD3 (Fujimoto et al., 2018) and IQL (Kostrikov et al., 2021) as the representative offline RL algorithms to obtain policies.
5.1 Benchmarking zero-shot performance for continuous control
Image-language reward Video-language reward RLZero
IQL TD3 (Base Model) TD3 IQL
Walker
Quadruped
Cheetah
Stickman
[…]
The ability to specify prompts and generate agent behavior allows us to explore complex behaviors that might have required complicated reward function design. We curate a set of 25 tasks across 4 DM-control environments. Each of the agents has unique capabilities as a result of its embodiment, and the prompts are specified to be reasonable tasks to expect for the specific domain. Furthermore, we filtered out prompts for which our off-the-shelf video generation model was unable to faithfully generate videos. We discuss this more in Appendix B.2. For each prompt, we generate behaviors for 5 seeds. The performance of any given method is evaluated as the win rate over the base method. We chose the base model for our comparisons as the policies trained via TD3 on image-language rewards. For each seed, we present the observation frames that the policy generated by different methods observes and pass it to a Multimodal LLM capable of video understanding, which is used as a judge. Since the number of tokens can get quite large with the long default horizon of the agent (1000 horizon), we subsample the videos by choosing every 8 frames and selecting the first 64 frames of size $256\times 256$. We observed this subsampling to retain temporal consistency and the effective horizon ($8\times 32=256$) to be long enough to demonstrate the task requested by the prompt.
Table 1 demonstrates the win rates by different methods when evaluated by GPT-4o-preview. We find that RLZero achieves a win rate of $83.2\%$ when compared to the best baseline which achieves a win rate of $51.2\%$. Figure 4 shows examples of behaviors output by RLZero on some of the prompts from our evaluation set. We also consider another metric for comparison – embedding similarity between a video of the generated behavior and the text. We use InternVideo2 to embed the videos and take the cosine similarity with the prompt used to generate the behavior. Table 1 also shows the results for this metric of comparison. Unfortunately, we observed that the similarity score is frequently higher even for behaviors that differ significantly from the prompt. This points to a limitation of using this metric for evaluation. Some reasons for this failure could be the limited context length of 8 for the video embedding model or a misalignment between video and text embedding vectors (Liang et al., 2022).
5.2 Can RLZero succeed at cross-embodiment imitation?
The intermediate stage in RLZero of matching the closest observations in the offline dataset to a frame from a video is based on semantic similarity. This means that we are not restricted to generating videos in the same domain of the agent and still expect semantic search to generalize for out-of-domain matching. Subsequently, we can skip the imagine step completely if we are given an expert demonstration. To investigate this, we consider a collection of videos scraped from Youtube as well as videos generated by open-source video generation tools like MetaAI and empirically test if RLZero is able to replicate the behaviors. We focus on the Stickman environment for our experiments here as it reflects human embodiment closely and allows us to use human videos from the internet.
@@ 5.2 Can RLZero succeed at cross-embodiment imitation?
Figure 5: Examples for cross embodied imitation: RLZero can mimic motions demonstrated in YouTube or AI generated videos zero-shot.
@@
Table 4 lists the set of 10 videos we use for cross-embodiment imitation. We use the win rate metric computed with GPT-4o with a prompt similar to Section 5.1, but modified to take in the frames from the original video instead of a specified task description. We compare against SMODICE (Ma et al., 2022a) which allows for using state-only observational data in conjunction with suboptimal offline data for imitation learning. This allows us to ablate the quality of imitation produced by a successor measure-based method that uses one model for all tasks as opposed to SMODICE that trains a new model for each task. RLZero achieves a win rate of 80% against SMODICE. This matches the observation from Pirotta et al. (2023) that DICE-based methods lag behind in performance on observation-only imitation tasks. Table 4 shows the win rate per task by RLZero across 10 youtube video imitation tasks. Figure 5 shows a qualitative comparison of the video and the obtained behavior on a few videos.
5.3 Ablation and Failure Cases
Imagination-free behavior generation: While the imagine, project, and imitate framework allows for interpretability into agent’s behavior, we investigate if we can amortize the imagination and embedding search cost by directly mapping the language embedding to the skill embedding in the Behavior Foundation Model’s latent space. For this, we consider sampling $z$ uniformly in the latent space of the BFM and embedding the generated image observation sequence through a ViFM, which we denote by $e$. Given the observation sequence, we generate the $z_{imit}$ using the zero-shot inference process and learn a mapping from $e\rightarrow z_{imit}$ using a small 3-layer MLP. On the same tasks considered in Table 5, we observe imagination-free RLZero to have a win rate of 65.71% over TD3 base model on Walker environment when compared to RLZero that had a win rate of $91.4\%$. A more thorough explanation of imagination-free RLZero can be found in Appendix B.6.

### v2

@@ 5 Experiments
Table 1: Winrates computed by GPT-4o of policies output by different methods when compared to a base policies trained by TD3+Image-language reward. Bolded distribution-matching returns denote statistically significant improvement over the second best method under a Mann-Whitney U test with a significance level of 0.05.
Setup: We consider continuous control tasks from the DeepMind control suite (Cheetah, Walker, Quadruped, Stickman) and HumEnv (3D Humanoid). The selected environments are diverse and complex, with their proprioceptive state space ranging from 17-358 dimensions and action space ranging from 6-69 dimensions. All of the environments use an environment horizon of 1000 steps, except HumEnv, which has a horizon of 300 steps. We use Cheetah, Walker, Quadruped, and Stickman environments to demonstrate language to behavior generation capabilities of RLZero and use Stickman and HumEnv environments to demonstrate cross-embodied abilities of RLZero to directly produce behaviors from out-of-distribution videos, such as those available on YouTube or AI-generated. Our choice of Stickman and HumEnv environments for cross-embodied behavior generation is motivated by the ease of availability of third-person individual human videos on internet as these environments reflect human morphology. We use an off-the-shelf pretrained BFM [74] trained for HumEnv with large MoCap datasets of real human behavior without assuming access to dataset used for training the BFM. This allows us to demonstrate the generality of the RLZero framework to plug a policy model trained with a different unsupervised RL algorithm to produce behaviors from cross-embodied videos. Our evaluation spans 25 tasks for language-to-behavior and 17 tasks for video-to-behavior.
Implementation: For the environments - Cheetah, Walker, Quadruped, and Stickman, we collect a dataset $d^{O}$ of environmental interactions of the form {$s,a,s^{\prime}$} purely by a random exploration algorithm RND [10] (10 million transitions for Stickman and 5 million transitions for the rest of the environments). In the Stickman environment, we additionally augment this dataset with replay buffers of the agent trained for walking and running to increase the diversity of behaviors, increasing the dataset size by another 2 million transitions. The detailed composition of the datasets can be found in Appendix B.3. We use the method from Section 4.1 to learn a text-conditioned video generation model using dataset $d^{O}$ without requiring any in-domain task-trajectory labeling. We use a successor-feature based unsupervised RL algorithm, Forward-Backward [76], to pretrain a general purpose policy BFM that can output a near-optimal policy given any reward function without test-time training or learning. For the projection step involved in language to behavior generation, we use SigLIP [89], an off-the-shelf image-text embedding model.
@@
Figure 3: Examples for cross embodied imitation: RLZero can mimic motions demonstrated in YouTube or AI generated videos zero-shot. Top 2 rows: Stickman (2D Humanoid), Bottom 2 rows: SMPL 3D Humanoid.
Baselines: For our evaluations, we consider the setting where the agent is not allowed environment interactions for learning during test time for a fair comparison and all baselines have access to same offline dataset $d^{O}$. This setting reflects the ability of the agents to use prior interaction data to learn meaningful behaviors. We compare against approaches that use foundation models trained with internet scale data to label rewards for offline transitions conditioned on text. These rewards can be computed using cosine similarity with text prompt per image observation of the agent as in VLM-RL [62] or per sequence of image observations using a video-embedding model. We use SigLIP as our image-text embedding model and InternVideo2 as a video-text embedding model for a fair comparison. Video-language embeddings take into account context and can potentially lead to more accurate reward estimation. Once the rewards are obtained, we use TD3 [23] and IQL [37] as the representative offline RL algorithms to obtain policies. We also compare to a recent approach GenRL, that performs model-based RL in the environment to learn a policy at test time by maximizing similarity to text-conditioned generated video. All baselines train a seperate policy for each task whereas RLZero trains a single policy for all tasks per environment using Forward-Backward unsupervised RL algorithm that uses TD3 as the underlying RL method.
Evaluation Protocol: We provide both qualitative and quantitative comparisons with baselines. The evaluation of behaviors generated given a text prompt or a video can be challenging as, unlike the traditional RL setting, we do not have access to a ground truth reward function. Instead, we have prompts that can be inherently ambiguous but reflect the reality of human-robot interaction. An obvious evaluation metric is to ask humans how much the generated behavior resembles their expectation of the behavior given the prompt. We use multimodal LLMs to qualitatively evaluate such preferences as a proxy to human preferences, as recent studies [14] have shown them to be correlated (up to 79.3%). To quantitatively compare the generated behaviors, we compute returns under a discriminator as a reward function that is trained to distinguish proprioceptive states corresponding to the projected frames of video (imagined or real) vs. offline dataset $d^{O}$ similar to [46].
5.1 Benchmarking Language to Continuous Control
The ability to specify prompts and generate agent behavior allows us to explore complex behaviors that might have required complicated reward function design. We curate a set of 25 tasks across 4 DM-control environments. Each of the agents has unique capabilities as a result of its embodiment, and the prompts are specified to be reasonable tasks to expect for the specific domain. For each prompt, we generate behaviors for 5 seeds. The qualitative performance of any given method is evaluated as the win rate over the base method. We chose the base model for our comparisons as the policies trained via TD3 on image-language rewards. For each seed, we present the observation frames that the output policy generates and pass it to a Multimodal LLM capable of video understanding, which is used as a judge. Since the number of tokens can get quite large with the long default horizon of the agent (1000 frames), we subsample the videos by choosing every 8 frames and selecting the first 64 frames of size $256\times 256$. We observed this subsampling to retain temporal consistency and the effective horizon ($8\times 32=256$) to be long enough to demonstrate the task requested by the prompt. We consider a more fine-grained quantitative metric for comparison – average return under the distribution-matching reward function. We learn a discriminator between the projected states for a given imagination ($\rho^{E}$) and the offline interaction dataset $\rho$. Under the shaped reward function (Section 4.3), $r(s)=e^{\nu}(s)=\rho^{E}/\rho$ we compute returns for all the methods; a higher return indicates that a method is better able to match the projected imaginations.
Video Descriptions SMODICE RLZero
Stickman (2D Humanoid) Human in backflip position — ($16.05\pm 3.44$) 2/5 ($14.27\pm 0.02$)
Downward-facing dog yoga pose — ($15.63\pm 2.74$) 1/5 ($14.11\pm 0.04$)
Cow yoga pose — ($6.72\pm 7.81$) 5/5 ($\mathbf{14.99\pm 0.02}$)
Downward dog one leg raised — ($7.04\pm 3.32$) 5/5 ($\mathbf{15.35\pm 0.03}$)
Lying on back, one leg raised — ($9.07\pm 2.07$) 5/5 ($\mathbf{12.55\pm 0.01}$)
Lying on back, both legs raised — ($8.45\pm 3.71$) 5/5 ($\mathbf{9.55\pm 0.06}$)
High plank yoga pose — ($12.83\pm 7.76$) 5/5 ($\mathbf{15.08\pm 0.03}$)
Sitting down, legs out front — ($12.01\pm 1.06$) 5/5 ($\mathbf{15.24\pm 0.02}$)
Warrior III pose — ($17.27\pm 5.09$) 3/5 ($15.22\pm 0.01$)
Front splits — ($13.44\pm 2.39$) 4/5 ($\mathbf{15.21\pm 0.01}$)
HumEnv (3D Humanoid) A karate kick position — ($50.02\pm 0.02$) 5/5 ($\mathbf{199.90\pm 0.01}$)
A cat doing a handstand — ($0.19\pm 0.24$) 5/5 ($\mathbf{199.86\pm 0.02}$)
An arabesque ballet position — ($10.02\pm 0.01$) 5/5 ($\mathbf{199.92\pm 0.04}$)
Animated wikiHow cartwheel — ($0.19\pm 0.24$) 5/5 ($\mathbf{199.91\pm 0.04}$)
Running — ($58.08\pm 0.46$) 5/5 ($\mathbf{199.87\pm 0.01}$)
Lying crunches — ($0.10\pm 0.20$) 5/5 ($\mathbf{199.89\pm 0.04}$)
Plank position — ($0.048\pm 0.03$) 5/5 ($\mathbf{199.91\pm 0.04}$)
Table 2: Cross Embodied Evaluation: Distribution Matching Return and Winrates
Table 5 demonstrates the win rates by different methods when evaluated by GPT-4o-preview as well as distribution matching returns. We find that RLZero achieves a win rate of $83.2\%$ when compared to the best baseline (GenRL), which achieves a win rate of $76.8\%$. RLZero shows statistically significant improvement over baselines under distribution matching returns. This indicates that RLZero is able to produce qualitatively better behaviors as well as do quantitatively better imitation than baselines. An important point to note is that all methods except RLZero requires test-time learning, whereas RLZero provides a gradient-free approach to instantly generating behaviors from text prompts. As an example, the closest performing baseline, GenRL requires test-time training with every task averaging $\approx$ 3-5 hours of training on NVIDIA-A40 GPU compared to our method, which requires $\approx$ 25 seconds to output a policy. The main bottleneck in RLZero is the projection step, which requires finding the closest real observation to each frame under a CLIP-like similarity metric and can be made significantly faster by using vector-databases specialized for retrieval. Figure 7 shows examples of behaviors output by RLZero on some of the prompts from our evaluation set. Additionally, all baselines require training a new policy model at test time for each prompt, as opposed to RLZero, which uses a single policy, making our approach parameter efficient.
5.2 Video to Continuous Control: Can RLZero succeed at cross-embodied imitation?
The intermediate stage in RLZero of matching the closest observations in the offline dataset to a frame from a video is based on semantic similarity. This means that RLZero can generalize to out-of-domain demonstrations building on the zero-shot generalization of CLIP-like representations. Subsequently, we can skip the imagine step completely if we are given an expert video demonstration. To investigate this, we consider a collection of videos scraped from Youtube as well as videos generated by open-source video generation tools like MetaAI and empirically test if RLZero is able to replicate the behaviors. We focus on Humanoid environments (Stickman, i.e 2D Humanoid, and HumEnv, i.e 3D Humanoid) here. For HumEnv [42], we use an open-source BFM [74] for policy generation given a reward function and do not assume access to the dataset used for training the BFM.
@@ 5.2 Video to Continuous Control: Can RLZero succeed at cross-embodied imitation?
Table 2 shows the results for cross-embodied imitation across 17 video-clips. We use similar metrics to Section 5.1, but modified the GPT-4o prompt to take in the frames from the original video instead of a specified task description. We compare against SMODICE [46] which allows for using state-only observational data in conjunction with suboptimal offline data for imitation learning. This allows us to ablate the quality of imitation produced by a successor measure-based method that uses one policy model for all tasks as opposed to SMODICE which trains a new policy for each task. GenRL requires a world model of the environment which is not available for HumEnv as we only have access to the pretrained policy. RLZero achieves a win rate of 80% against SMODICE for Stickman and 100% for HumEnv. This matches the observation from [57] that DICE-based methods lag behind in performance on observation-only imitation tasks. Figure 3 shows a qualitative comparison of the video and the obtained behavior on a few videos. Details about videos used can be found in Appendix C.3.
5.3 Ablation and Limitations